OpenStudy (anonymous):

the exact solution of (du/dt)+c(du/dx)=v(d^2u/dx^2) where d is for partial derivative

OpenStudy (nikvist):

Do you think \[\frac{\partial u}{\partial t}+c\frac{\partial u}{\partial x}=v\frac{\partial^2 u}{\partial x^2}\] c,v are constants